# Singularity-free cosmology from interactions in the dark sector

### v1

@@ Abstract
We study the dynamics of Friedmann-Lemaître-Robertson-Walker models where a dark energy component with a quadratic equation of state (EoS) nonlinearly interacts with cold dark matter. Thus, two energy scales naturally come into play: $\rho_{*}$ is the scale at which the nonlinearity of the EoS becomes relevant; $\rho_{i}$ is the energy scale around which the interaction starts to play an important role in the dynamics. Our focus is to understand whether there are parameter ranges for this system that can produce non-singular bouncing and emergent cosmologies for any initial condition. We complete a dynamical systems analysis, and find the parameter range such that trajectories always expand from a high energy non-singular de Sitter state. For flat and negative curvature models this de Sitter state is represented by a fixed point, the asymptotic past from which the universe emerges from. We find a subset of positive curvature models that during contraction get arbitrarily close to the de Sitter state, thus having a quasi-de Sitter bounce, then emerge from the bounce and expand, evolving toward spatial flatness. We find that the dimensionless parameter $q\equiv\rho_{*}/\rho_{i}$, which measures the relative strength of the nonlinear terms in the system, plays a crucial role in the topology of the phase space. When $q<3$, some trajectories expand toward a singularity, while others evolve toward a low energy cosmological constant at late-times, with a subset going through a decelerated matter dominate era before the final acceleration. When $q>3$, all trajectories are non-singular, and evolve toward a late-time cosmological constant. We find a subclass in this case in which all trajectories have at least one decelerated matter dominated phase, and accelerate at late-times. Therefore, the nonlinear interacting cosmology presented here allows for a subclass of models that are singularity-free and qualitatively realistic for any initial condition.
@@ I Introduction
The $\Lambda$CDM model plus inflation is our current Standard Model of Cosmology, as it has been the most consistent theory, agreeing with a wealth of observations Aghanim et al. (2020a, b). However, there are theoretical problems which require addressing. One problem is that of singularities. Assuming the strong energy conditions (SEC) holds, cosmological singularities are inevitable in General Relativity (GR), as shown by Hawking and Penrose Penrose (1965); Hawking (1965, 1966a, 1966b, 1967); Hawking and Penrose (1970); Hawking and Ellis (1973). However, their current interpretation is that they represent points where GR breaks down Joshi (1993); Hawking and Israel (1979); Gibbons et al. (2003); Ashtekar et al. (2015). There are two ways we can tackle this problem: either we can develop a quantum theory of gravity to replace GR at Planck energies and above, or we can avoid a singularity by developing a non-singular classical framework for high but sub-Planckian energies. One alternative to an initial singularity are bouncing cosmologies, which transition from an initially contracting period to an expanding phase Gasperini and Veneziano (1993); Khoury et al. (2001); Kallosh et al. (2001); Steinhardt and Turok (2002); Novello and Bergliaffa (2008); Wands (2008); Maier et al. (2012, 2013); Battefeld and Peter (2014); Cai et al. (2017); Brandenberger and Peter (2017); Cai and Piao (2017); Ijjas and Steinhardt (2018, 2019); Ganguly and Bruni (2019); Ilyas et al. (2020); Battista (2021); Zhu et al. (2021). An emergent universe is another alternative, originally conceived as initially quasi-static, expanding from a non-singular Einstein universe state with positive curvature Ellis and Maartens (2003); Ellis et al. (2004); Mukherjee et al. (2006); Verlinde (2017); Ellis et al. (2016); Parisi et al. (2007); Mulryne et al. (2005); Bonanno et al. (2017); Gionti (2017); Sengupta (2024); Barrow et al. (2003). The problem with this scenario is that it is not self-consistent, at least classically, as it relies on some previously existing mechanism to establish such very special initial conditions, which otherwise is a set of measure zero in phase space Burkmar and Bruni (2024). In Burkmar and Bruni (2024) we contended that a different emergent scenario can be built, where all trajectories in phase space representing expanding models evolve from a past non-singular de Sitter phase.
@@
Another problem facing the Standard Model of Cosmology is that of dark energy. The cosmological constant $\Lambda$ is the simplest form of dark energy that we have to explain the accelerated expansion of our Universe Riess et al. (1998); Perlmutter et al. (1998). However, the observed value of $\Lambda$ is at odds by up to 120 orders of magnitude with theoretical estimates of the contributions to the effective cosmological constant expected from quantum field theory (QFT) Weinberg (1989); Straumann (1999); Ellis et al. (2012).
It has also been debated as to whether $\Lambda$ is a true constant, or an approximation to a more complex dynamic dark energy. Beyond a cosmological constant, an interacting vacuum has been widely studied Wands et al. (2012); De-Santiago et al. (2012); Hogg and Bruni (2021); Kaeonikhom et al. (2020); Quercellini et al. (2008), where energy-momentum is transferred between the vacuum and other matter fields. Interactions between the dark sector can alleviate the coincidence problem: that is, the energy densities of cold dark matter and dark energy are of the same order of magnitude. An interacting vacuum model has also been studied in the context of non-singular cosmologies Bruni et al. (2022). Other dynamic dark energy models have been studied where there are no interactions in the dark sector, but where the dark energy has a non-linear quadratic equation of state (EoS) Ananda and Bruni (2006a, b). A quadratic EoS is the simplest non-linear EoS one can study, and can give rise to non-singular bouncing and emergent cosmologies.
In our previous work, which we will refer to as Paper I Burkmar and Bruni (2023), we extended Ananda and Bruni (2006a, b), and studied a dark energy with a non-linear quadratic EoS. We included dark matter and radiation in the set-up, without any interaction term between the components. Non-singular models were possible, however when we quantitatively set the energy scales for the dark energy, we lost a decelerated period where large scale structure could form. In this paper, we want to understand whether including an interaction between the dark components would allow for qualitatively realistic non-singular bouncing and emergent models, which could be consistent with observations, with quantitatively realistic energy scales. In particular, we would like to understand if there are parameter ranges for which these non-singular models are general for any initial condition. The analysis for this work is available through GitHub, provided the reader has a Mathematica license Burkmar (2025).
@@
This paper is organized as follows. In Sec. II, we present the system of equations for the dark energy, dark matter and the Hubble expansion function in terms of dimensionless variables. In Sec. III, we explore the parameter space, and set parameters such that the topology of the phase space is qualitatively meaningful. With this, here and in the following, we mean that expanding models go through a matter-dominated decelerated phase before the late-time acceleration of the expansion, thereby - at least qualitatively - corresponding to the observed universe. In Sec. IV, we study the dynamics of the phase space when $q$, which measures the relative strength of the nonlinear terms in the system, satisfies the condition $q<3$, and in Sec. V we study the dynamics when $q>3$, showing that in this case all universe models are singularity-free. We present our conclusions in Sec. VI. In this paper, we employ natural units such that $18\pi G=c=$ 1.
@@ II.1 Physical variables
The evolution of Friedmann-Lemaître-Robertson-Walker (FLRW) models is described by a system of ODEs, which consist of the energy conservation equations for each component, as well as the Raychaudhuri equation which describes the expansion scalar $H$, also known as the Hubble function. The Friedmann equation (the Hamiltonian constraints) relates $H$ to the energy densities of the components and to the 3-curvature of the constant time hypersurfaces.
In this paper we consider dark matter and dark energy with a non-gravitational interaction term. We take dark matter to be cold (CDM) and pressureless. Then, its energy conservation equation is expressed as a sum of the standard matter part and the interaction,
@@
where $\rho_{m}$ is the dark matter energy density, $\rho_{x}$ is the dark energy density, and $\rho_{i}$ is the energy scale characterizing the interaction; overdots are derivatives with respect to time $t$. The dark energy has a self-interaction term with a quadratic EoS, and the pressure of the dark energy, $P_{x}$, is expressed as  [^4] The parametrization here is different from Paper I Burkmar and Bruni (2023).
@@
where $w_{x}$ is its equation of state parameter, $\epsilon$ sets the sign of the quadratic term, $\rho_{\Lambda}$ is the effective low-energy cosmological constant, which as we will see in Sec. III is the low energy attractor, and $\rho_{*}$ is the characteristic energy scale of the dark energy. The energy conservation equation for the dark energy then becomes
@@
The interaction term has been set such that if the initial conditions for the dark energy density satisfy $\rho_{x}>0$ and $\dot{\rho_{x}}<0$, then the dark energy density is always positive. This also implies that during expansion the energy flows from the dark energy to the dark matter. Depending on the values of parameters, it is possible for the dark energy to violate the null energy condition (NEC) during certain periods of its evolution. Therefore, the dark energy can be phantom, meaning $\rho_{x}$ increases with expansion during those periods. For $\rho_{m}=0$, it is clear from (3) that $\rho_{x}=\rho_{\Lambda}$ represents a cosmological constant: as we show later, this is the low-energy attractor of our dynamical system. We think of $\rho_{\Lambda}$ as being close to the observed dark energy density. Similarly, for $\rho_{m}=0$ the energy scale $\rho_{*}$ also gives a cosmological constant, albeit a high energy and unstable one. We assume that $\rho_{*}$ and $\rho_{i}$ are energy scales between the Planck scale and that of inflation, such that at lower energies the dynamics is similar to that of the standard model of cosmology.
@@
is the related acceleration equation. Eq. (4) admits a first integral, the Friedmann equation:
@@ II.2 Dimensionless variables
We normalize the above system of equations with respect to the characteristic energy scale of the dark energy, $\rho_{*}$, thereby introducing dimensionless variables and parameters. Following Paper I Burkmar and Bruni (2023) and Ananda and Bruni (2006a, b), we define these variables as:
@@
The variables $x$ and $z$ are the normalized dark energy density and the normalized dark matter energy density, respectively, $y$ is the normalized Hubble expansion function, and $\eta$ is the normalized time variable. In Paper I Burkmar and Bruni (2023), we included radiation in the set-up, which we do not do here, and the system of ODEs in Paper I could be solved analytically, which is not possible here. This means unlike in Paper I, we cannot project the dynamics onto lower-dimensional phase spaces. In any case, including radiation would only give a transition regime between the high energies, dominated by the dark energy and the interaction, and the matter dominated era. We also introduce the dimensionless parameters,
@@
where $\mathcal{R}$ is the ratio of the low energy effective cosmological constant $\rho_{\Lambda}$ to the characteristic energy scale of the dark energy $\rho_{*}$, which takes a value in the range $10<\mathcal{R}<$ 1; $q$ is the strength of interaction parameter. Our system of equations then becomes:
@@
where primes indicate differentiation with respect to the time variable $\eta$. The acceleration equation is then
@@
where $y$ can be connected to the scale factor $a$ through the expression $y=a^{\prime}/a$. We only consider the region of phase space where $x,z>0$, such that the energy densities of the dark energy and dark matter are always positive. From Eq. 10, it is clear that $z=0$ is an invariant manifold that cannot be crossed. As we will show later, although $x=0$ is not an invariant manifold, initial conditions with $x>0$ and $\dot{x}<0$, such that $x>0$ always in their past, can never cross $x=0$ from above.
@@ III Analysis of Parameters
III.1 2-D phase space
For this system to be non-singular, and to be compatible with observations, we require all trajectories to emerge from a high energy repellor during expansion, and tend toward an attractor at low energy. Therefore, we need to understand the parameter space that allows for this topology. To do this, we need to look at the 2-D phase space. Taking $y>0$ ($H>0$), and using the relation
@@
we now change the time variable to $\ln(a)$. Our equations in 2-D become
@@
where now primes indicate differentiation with respect to $\ln({a})$. We then need to find the fixed points for the 2-D system, which are shown in Table 1, and their eigenvalues, which can be found in Table 2. We classify all the fixed points in the above system as representing de Sitter models, because they have constant energy densities $x^{\prime}=z^{\prime}=0$, where $y\neq 0$ and can vary for positively or negatively curved de Sitter models. In general, this system admits three fixed points. The fixed point $dS_{1+}$ comes from the interaction term, and $dS_{2+}$ and $dS_{3+}$ come from the non-linear equation of state for the dark energy in Eq. (2).
@@
$dS_{2\pm}$ 00 $\mathcal{R}$
$dS_{3\pm}$ 00 $-\frac{(1+w_{x})}{\epsilon}$
Table 1: The fixed points of the $z$-$x$ system. $dS_{+}$ denotes an expanding (+) de-Sitter universe ($x^{\prime}=z^{\prime}=0$) when $y>0$, and $dS_{-}$ denotes a contracting de-Sitter universe when $y<0$. In the 2-D phase spaces, only the $dS_{+}$ fixed points are plotted as we have fixed $y>0$ so the trajectories are expanding.
We want to find parameters such that when $y>0$ ($H>0$), the high energy fixed point $dS_{1+}$ is a repellor, so trajectories emerge from a non-singular de Sitter fixed point during expansion. We also require that the low energy fixed point $dS_{2+}$ is an attractor, such that trajectories asymptotically approach a low energy cosmological constant to ensure the dynamics qualitatively matches our observed Universe. To understand whether this system allows for a high energy repellor and low energy attractor, we need to look at the eigenvalues of the fixed points, the sign of which depends on the specific values of $\epsilon$, $w_{x}$ and $q$. For $dS_{1+}$ to be a repellor, its eigenvalues must have positive real parts, and for $dS_{2+}$ to be an attractor its eigenvalues must have negative real parts. Both of these fixed points must also exist at $x>0$ and $z>0$, so that the energy densities of the dark energy and dark matter are positive. In the following subsections, we analyze different parameter ranges to understand whether $dS_{1+}$ and $dS_{2+}$ can have the stability character we require.
@@
$dS_{1+}$ $-\frac{q^{2}\mathcal{R}(1+w_{x})+9\epsilon+\sqrt{-12q(q+qw_{x}+3\epsilon)(q\mathcal{R}-3)+\left[q^{2}\mathcal{R}(1+w_{x})+9\epsilon\right]^{2}}}{2q}$ $-\frac{q^{2}\mathcal{R}(1+w_{x})+9\epsilon-\sqrt{-12q(q+qw_{x}+3\epsilon)(q\mathcal{R}-3)+\left[q^{2}\mathcal{R}(1+w_{x})+9\epsilon\right]^{2}}}{2q}$
$dS_{2+}$ $-3+q\mathcal{R}$ $-3(1+w_{x}+\epsilon\mathcal{R})$
$dS_{3+}$ $-3-\frac{q(1+w_{x})}{\epsilon}$ $3(1+w_{x}+\epsilon\mathcal{R})$
Table 2: The eigenvalues of the fixed points of the $z$-$x$ system. $dS_{+}$ denotes an expanding (+) de-Sitter universe ($x^{\prime}=z^{\prime}=0$). Note that for the contracting de-Sitter fixed points $dS_{-}$, the eigenvalues change sign.

### v2

@@ Abstract
We study the dynamics of Friedmann-Lemaître-Robertson-Walker models where a dark energy component with a quadratic equation of state (EoS) nonlinearly interacts with cold dark matter. Thus, two energy scales naturally come into play: $\rho_{*}$ is the scale at which the nonlinearity of the EoS becomes relevant; $\rho_{i}$ is the energy scale around which the interaction starts to play an important role in the dynamics. Our focus is to understand whether there are parameter ranges for this system that can produce nonsingular bouncing and emergent cosmologies for any initial condition. We complete a dynamical systems analysis, and find the parameter range such that trajectories always expand from a high energy nonsingular de Sitter state. For flat and negative curvature models this de Sitter state is represented by a fixed point, the asymptotic past from which the Universe emerges. We find a subset of positive curvature models that during contraction get arbitrarily close to the de Sitter state, thus having a quasi-de Sitter bounce, then emerge from the bounce and expand, evolving toward spatial flatness. We find that the dimensionless parameter $q\equiv\rho_{*}/\rho_{i}$, which measures the relative strength of the nonlinear terms in the system, plays a crucial role in the topology of the phase space. When $q<3$, some trajectories expand toward a singularity, while others evolve toward a low energy cosmological constant at late times, with a subset going through a decelerated matter dominate era before the final acceleration. When $q>3$, all trajectories are nonsingular, and evolve toward a late-time cosmological constant. We find a subclass in this case in which all trajectories have at least one decelerated matter dominated phase, and accelerate at late times. Therefore, the nonlinear interacting cosmology presented here allows for a subclass of models that are singularity-free and qualitatively realistic for any initial condition.
@@ I Introduction
[…]
@@
Another problem facing the standard model of cosmology is that of dark energy. The cosmological constant $\Lambda$ is the simplest form of dark energy that we have to explain the accelerated expansion of our Universe Riess et al. (1998); Perlmutter et al. (1998). However, the observed value of $\Lambda$ is at odds by up to 120 orders of magnitude with theoretical estimates of the contributions to the effective cosmological constant expected from QFT Weinberg (1989); Straumann (1999); Ellis et al. (2012). Various attempts have been made to resolve this problem within the Friedmann-Lemaître-Robertson-Walker (FLRW) model; for example, certain running vacuum models in QFT may be able to alleviate the fine-tuning, see e.g. Ref. Solà Peracaula (2022) and references therein.
[…]
In our previous work, which we will refer to as Paper I Burkmar and Bruni (2023), we extended Refs. Ananda and Bruni (2006a, b), and studied a dark energy with a nonlinear quadratic EoS. This dark energy was past asymptotic to a high energy effective cosmological constant during expansion, and evolved towards a low energy effective cosmological constant at late times. We included dark matter and radiation in the setup, without any interaction term between the components. Nonsingular models were possible; however, when we quantitatively set the energy scales for the dark energy, such that the high energy effective cosmological constant was between Planck and inflationary energy scales, and the low energy effective cosmological constant was close to the observed value of $\Lambda$, we lost a decelerated period where large scale structure could form. In this paper, we want to understand whether including an interaction between the dark components would allow for qualitatively realistic nonsingular bouncing and emergent models, which could be consistent with observations, with quantitatively realistic energy scales. We introduce an interaction term which is quadratic and proportional to the product of the dark matter and dark energy densities, such that only the high energy dynamics are modified where we want to avoid singularities. We expect a realistic model to bounce or emerge between Planck and inflationary scales, and have a low energy, dark energy dominated acceleration phase at late times. At high energy, we would need to constrain the model with current bounds on primordial gravitational waves. If we linearize our equations, we obtain a low energy model which has already been tested against, and is consistent with data from the Wilkinson Microwave Anisotropy Probe (WMAP), the Sloan Digital Sky Survey (SDSS) and the Hubble Space Telescope Balbi et al. (2007); Pietrobon et al. (2008). In addition, models with a linear interaction term have been tested against supernovae data Quercellini et al. (2008), therefore we expect the late-time behavior of our model to be consistent with these observations. In this paper, we would like to understand if there are parameter ranges in which nonsingular models emerge between Planck and inflationary scales, have a decelerated expansion phase, and asymptotically tend towards a low energy cosmological constant for any initial condition. We note that we are really interested in the qualitative behavior of the model we present in this paper. Here our aim is to understand whether a singularity can be avoided at high energy, with late-time behavior that is qualitatively approximate to the observed universe, such that we obtain a low energy, accelerating phase dominated by dark energy. However, in light of the DESI results Abdul-Karim et al. (2025), we appreciate that the late-time behavior of the model we consider may need modifying in order to accommodate a dark energy that is dynamical at low energy, rather than asymptotically constant. The analysis for this work is available through GitHub, provided the reader has a Mathematica license Burkmar (2025).
@@
This paper is organized as follows: in Sec. II, we present the system of equations for the dark energy, dark matter and the Hubble expansion function in terms of dimensionless variables. In Sec. III, we explore the parameter space, and set parameters such that the topology of the phase space is qualitatively meaningful. With this, here and in the following, we mean that expanding models go through a matter dominated decelerated phase before the late-time acceleration of the expansion, thereby – at least qualitatively – corresponding to the observed universe. In Sec. IV, we study the dynamics of the phase space when $q$, which measures the relative strength of the nonlinear terms in the system, satisfies the condition $q<3$, and in Sec. V we study the dynamics when $q>3$, showing that in this case all universe models are singularity-free. We present our conclusions in Sec. VI. In this paper, we employ natural units such that $8\pi G=c=1$.
@@ II.1 Physical variables
The evolution of FLRW models is described by a system of ordinary differential equations (ODEs), which consist of the energy conservation equations for each component, as well as the Raychaudhuri equation which describes the expansion scalar $H$, also known as the Hubble function. The Friedmann equation (the Hamiltonian constraints) relates $H$ to the energy densities of the components and to the three-curvature of the constant time hypersurfaces.
In this paper we consider dark matter and dark energy with a nongravitational interaction term. We take dark matter to be cold (CDM) and pressureless. Then, its energy conservation equation is expressed as a sum of the standard matter part and the interaction,
@@
where $\rho_{m}$ is the dark matter energy density, $\rho_{x}$ is the dark energy density, and $\rho_{i}$ is the energy scale characterizing the interaction, i.e. the scale at which the interaction term becomes significant in the dynamics; overdots are derivatives with respect to time $t$. An interaction term which is proportional to the product of the energy densities of the dark components has been studied before, and has been found to help alleviate the coincidence problem Lip (2010). This type of model can also admit solutions where the density parameters oscillate, such that the dark energy and dark matter repeatedly switch between being the dominant component, and their energy densities do not become infinite Arévalo et al. (2012); Perez et al. (2014); Aydiner (2018). In this paper, we implement a quadratic interaction term with the aim of only modifying the dynamics at high energy, to understand if the energy densities of the dark components remain finite, such that we can avoid singularities.
The dark energy has a self-interaction term with a quadratic EoS, and the pressure of the dark energy, $P_{x}$, is expressed as  [^4] The parametrization here is different from Paper I Burkmar and Bruni (2023).
@@
where $w_{x}$ sets its equation of state at low energy; $\epsilon$ sets the sign of the quadratic term; $\rho_{\Lambda}$ is the effective low energy cosmological constant, which as we will see in Sec. III is the low energy attractor; and $\rho_{*}$ is the characteristic energy scale of the dark energy, i.e. the scale at which the nonlinear terms in Eq. (2) become significant. We note that although this model is phenomenological, it has been shown that a perfect fluid with affine barotropic EoS can represent a standard or k-essence scalar field Quercellini et al. (2007). The energy conservation equation for the dark energy then becomes
@@
The interaction term has been set such that if the initial conditions for the dark energy density satisfy $\rho_{x}>0$ and $\dot{\rho_{x}}<0$, then the dark energy density is always positive. This also implies that during expansion the energy flows from the dark energy to the dark matter. Depending on the values of parameters, it is possible for the dark energy to violate the null energy condition (NEC) during certain periods of its evolution. Therefore, the dark energy can be phantom, meaning $\rho_{x}$ increases with expansion during those periods. For $\rho_{m}=0$, it is clear from Eq. (3) that $\rho_{x}=\rho_{\Lambda}$ represents a cosmological constant: as we show later, this is the low energy attractor of our dynamical system. We think of $\rho_{\Lambda}$ as being close to the observed dark energy density. Similarly, for $\rho_{m}=0$ the energy scale $\rho_{*}$ also gives a cosmological constant, albeit a high energy and unstable one. We assume that $\rho_{*}$ and $\rho_{i}$ are energy scales between the Planck scale and that of inflation. This is so our model does not interfere with particle production at lower energies, and so a decelerated matter dominated phase can follow. We want to understand if nonsingular models which admit a matter dominated decelerated phase are possible, particularly when we set realistic energy scales.
@@
is the related acceleration equation. Equation (4) admits a first integral, the Friedmann equation:
@@ II.2 Dimensionless variables
We normalize the above system of equations with respect to the characteristic energy scale of the dark energy, $\rho_{*}$, thereby introducing dimensionless variables and parameters. Following Paper I Burkmar and Bruni (2023) and Refs. Ananda and Bruni (2006a, b), we define these variables as
@@
The variables $x$ and $z$ are the normalized dark energy density and the normalized dark matter energy density, respectively, $y$ is the normalized Hubble expansion function, and $\eta$ is the normalized time variable. In Paper I Burkmar and Bruni (2023), we included radiation in the setup, which we do not do here, and the system of ODEs in Paper I could be solved analytically, which is not possible here. This means that unlike in Paper I, we cannot project the dynamics onto lower-dimensional phase spaces. In any case, including radiation would only give a transition regime between the high energies, dominated by the dark energy and the interaction, and the matter dominated era. We also introduce the dimensionless parameters,
@@
where $\mathcal{R}$ is the ratio of the low energy effective cosmological constant $\rho_{\Lambda}$ to the characteristic energy scale of the dark energy $\rho_{*}$, which takes a value in the range $0<\mathcal{R}<1$, and $q$ is the strength of the interaction parameter. We assume that $\rho_{\Lambda}$ is close to the observed value of $\Lambda$, and $\rho_{*}$ is between Planck and inflationary energy scales, such that the evolution is completely classical without the need of a quantum gravity era at high energy. Observations show that the current vacuum energy density is $\rho_{obs}\sim 10^{-47}$GeV4 Aghanim et al. (2020b); Prat et al. (2022). The Planck scale is approximately 120 orders of magnitude larger than $\rho_{obs}$, and if measurements of the tensor-to-scalar ratio fall within the range $0.001<r<0.1$, then it would point to a scale of inflation of $\rho_{{}_{inf}}\sim 10^{64}$GeV4 Bass (2015). These measurements provide a realistic range of $\mathcal{R}$ to be $10^{-120}<\mathcal{R}<10^{-111}$. In Sec. IV, we will use a larger value of $\mathcal{R}$ purely for the purpose of producing readable phase spaces. In terms of dimensionless variables, Eq. (2) becomes
$\frac{P_{x}}{\rho_{*}}=-\mathcal{R}(1+w_{x})+x(w_{x}-\epsilon\mathcal{R})+\epsilon x^{2}\,$, (9)
and our system of equations in Eqs. (3), (1) and (4) become
@@
where primes indicate differentiation with respect to the time variable $\eta$. We note that models with affine EoS $P=P_{0}+\alpha\rho$ have already been tested against, and were found to be consistent with, observational data from WMAP, SDSS and the Hubble Space Telescope Balbi et al. (2007); Pietrobon et al. (2008). If we linearize the EoS in Eq. (9), we obtain $P_{x}/\rho_{*}=-\mathcal{R}(1+w_{x})+x\left(w_{x}-\epsilon\mathcal{R}\right)$, and as the interaction term is quadratic, we can neglect it in Eq. (10) at late times. Therefore, the effective dark energy that we observe at low energy depends on $w_{x}$ and $\epsilon$, as $\mathcal{R}$ is already constrained to $10^{-120}<\mathcal{R}<10^{-111}$ in the model. We therefore expect that if we set the sign of $\epsilon$ and constrain $w_{x}$ with WMAP, SDSS and Hubble Space Telescope data, our model will fit observations, just as $\alpha$ was constrained in models with EoS $P=P_{0}+\alpha\rho$ which were found to be consistent with data Balbi et al. (2007); Pietrobon et al. (2008).
In dimensionless variables, the acceleration equation is
@@
where $y$ can be connected to the scale factor $a$ through the expression $y=a^{\prime}/a$. We only consider the region of phase space where $x,z>0$, such that the energy densities of the dark energy and dark matter are always positive. From Eq. (11), it is clear that $z=0$ is an invariant manifold that cannot be crossed. Although $x=0$ is not an invariant manifold, from Eq. (10) we see that along $x=0$, $x^{\prime}>0$; although phantom trajectories ($x^{\prime}>0$) can cross from below $x=0$, trajectories cannot cross $x=0$ from above. Therefore, as long as initial conditions are chosen satisfying $x>0$ and $\dot{x}<0$, such that we are in nonphantom region, this is sufficient for trajectories to always satisfy $x>0$, even if they become phantom at late times.
@@ III Analysis of Parameters
III.1 Two-dimensional phase space
For this system to be nonsingular, and to be compatible with observations, we require all trajectories to emerge from a high energy repellor during expansion, and tend toward an attractor at low energy. Therefore, we need to understand the parameter space that allows for this topology. To do this, we need to look at the two-dimensional phase space. Taking $y>0$ ($H>0$), and using the relation
@@ III.1 Two-dimensional phase space
we now change the time variable to $\ln(a)$. Our equations in two dimensions become
@@
where now primes indicate differentiation with respect to $\ln({a})$. We then need to find the fixed points for the two-dimensional system, which are shown in Table 1, and their eigenvalues, which can be found in Table 2. We classify all the fixed points in the above system as representing de Sitter models, because they have constant energy densities $x^{\prime}=z^{\prime}=0$, where $y\neq 0$ and can vary for positively or negatively curved de Sitter models. In general, this system admits three fixed points. The fixed point $dS_{1+}$ comes from the interaction term, and $dS_{2+}$ and $dS_{3+}$ come from the nonlinear equation of state for the dark energy in Eq. (2).
@@
$dS_{2\pm}$ 0 $\mathcal{R}$
$dS_{3\pm}$ 0 $\frac{(1+w_{x})}{\epsilon}$
Table 1: Fixed points of the $z$-$x$ system. $dS_{+}$ denotes an expanding (+) de Sitter universe ($x^{\prime}=z^{\prime}=0$) when $y>0$, and $dS_{-}$ denotes a contracting de Sitter universe when $y<0$. In the two-dimensional phase spaces, only the $dS_{+}$ fixed points are plotted as we have fixed $y>0$ so the trajectories are expanding.
We want to find parameters such that when $y>0$ ($H>0$), the high energy fixed point $dS_{1+}$ is a repellor, so trajectories emerge from a nonsingular de Sitter fixed point during expansion. We also require that the low energy fixed point $dS_{2+}$ is an attractor, such that trajectories asymptotically approach a low energy cosmological constant to ensure the dynamics qualitatively matches our observed Universe. To understand whether this system allows for a high energy repellor and low energy attractor, we need to look at the eigenvalues of the fixed points, the sign of which depends on the specific values of $\epsilon$, $w_{x}$ and $q$. For $dS_{1+}$ to be a repellor, its eigenvalues must have positive real parts, and for $dS_{2+}$ to be an attractor its eigenvalues must have negative real parts. Both of these fixed points must also exist at $x>0$ and $z>0$, so that the energy densities of the dark energy and dark matter are positive. In the following subsections, we analyze different parameter ranges to understand whether $dS_{1+}$ and $dS_{2+}$ can have the stability character we require.
@@
$dS_{1+}$ $\frac{q^{2}\mathcal{R}(1+w_{x})+9\epsilon+\sqrt{-12q(q+qw_{x}+3\epsilon)(q\mathcal{R}-3)+\left[q^{2}\mathcal{R}(1+w_{x})+9\epsilon\right]^{2}}}{2q}$ $\frac{q^{2}\mathcal{R}(1+w_{x})+9\epsilon-\sqrt{-12q(q+qw_{x}+3\epsilon)(q\mathcal{R}-3)+\left[q^{2}\mathcal{R}(1+w_{x})+9\epsilon\right]^{2}}}{2q}$
$dS_{2+}$ $3+q\mathcal{R}$ $3(1+w_{x}+\epsilon\mathcal{R})$
$dS_{3+}$ $3-\frac{q(1+w_{x})}{\epsilon}$ $3(1+w_{x}+\epsilon\mathcal{R})$
Table 2: Eigenvalues of the fixed points of the $z$-$x$ system. $dS_{+}$ denotes an expanding (+) de Sitter universe ($x^{\prime}=z^{\prime}=0$). Note that for the contracting de Sitter fixed points $dS_{-}$, the eigenvalues change sign.